Fermi National Laboratory
 
Operation Summary for March 23, 2012 - March 26, 2012
 

NuMI reported receiving 8.85E18 protons on target for the period from 3/19/12 to 3/26/12.

Friday March 23

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest experiments, and the Meson FTBFT-1017 experiment taking beam.

During the day shift, at 9:32 AM, Operations prepared Booster for an access to repair a Booster RF station (BRF3) flow meter; Operators loaded an energy conservation timeline into the Time Line Generator (TLG). Booster came out of access at 10:36 AM. Operations reestablished beam to NuMI and MiniBooNE at 10:55 AM. A Muon Ring person began a muon study using the antiproton target at 11:04 AM. The NuMI beam permit tripped off at 11:14 AM due to a power supply (V118) drifting low; everything reset. Operations established beam to T-1017 at 11:30 AM. At 12:22 PM, the H- Source tripped off and would not reset; Linac technicians investigated and had the source up by 12:30 PM. Operations reported that they had received the Operational Readiness Clearance (ORC) for the FTBF experiment T-1012 and they began taking parasitic beam behind the T-1017 experiment. MTA also took beam during this shift. SeaQuest and T-1017 made accesses during this shift.

During the evening shift, Operations disabled beam to MTA at 5:17 PM. Muon Ring studies ended at 6:16 PM. At 8:05 PM, a Linac RF station (LRF3) tripped off for the first of three times in five minutes due to a low driver output; it reset. A Linac person came in, made adjustments to the station, and had it running normally by 8:42 PM. SeaQuest made an access during this shift.

Saturday March 24

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, T-1017, and T-1012.

During the midnight shift, LRF3 tripped off at 12:11 AM; Operators contacted a Linac person who told Operators how to adjust the station to bring it back online. Linac resumed sending beam at 12:16 AM. At 1:39 AM, the MiniBooNE beam permit tripped off for the first of three times in five hours due to V118; it reset. SeaQuest made an access during this shift.

During the day shift, at 12:14 PM, LRF3 tripped off for the first of four times in about 90 minutes; it reset. A Linac person advised Operators on tuning the station.

During the evening shift, at 10:39 PM, the NuMI beam permit tripped off due to a bad beam position; it reset. At 10:52 PM, the NuMI beam permit tripped off due to V118; everything reset.

Sunday March 25

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, T-1017, and T-1012.

During the midnight shift, at 2:35 AM, a MI kicker (KPS4E) tripped off; it reset. At 6:48 AM, the NuMI beam permit tripped off due to a bad beam position; it reset. The MiniBooNE beam permit tripped off at 4:38 AM due to bad beam positions; it reset.

During the day shift, at 10:53 AM, the MI-20 LCW pump tripped off and would not reset; Operators contacted a MI person who said the machine could run without the pump until Monday. The NuMI beam permit tripped off for the first of two times at 12:28 PM due to V118; everything reset. The H- Source high voltage tripped off at 3:23 PM; it reset. T-1017 made an access during this shift.

During the evening shift, the NuMI beam permit tripped off at 5:21 PM due to V118; it reset. At 8:56 PM, the MiniBooNE beam permit tripped off due to a bad beam position; it reset.

Monday March 26

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, T-1017, and T-1012.

During the midnight shift, at 12:35 AM, the NuMI beam permit tripped off for the first of two times in about four and half hours; it reset. At 12:44 AM, the MiniBooNE beam permit tripped off for the first of two times in about one hour due to bad beam positions; it reset.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with the plans of Booster and MI personnel to conduct studies. SeaQuest continues to commission their beam line.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/26/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory